OpenStudy (anonymous):

e^2x= e((x+log5)/loge)-6

terenzreignz (terenzreignz):

lol... don't you just love how they complicate the simple? ^_^

terenzreignz (terenzreignz):

\[\Large e^{2x}= e^{\frac{x+\log 5}{\log e}}-6 \] Is this it?

OpenStudy (anonymous):

yes!

terenzreignz (terenzreignz):

Well this is a bit hard, isn't it? Perhaps we can use the change of base formula somewhere... do you know the change of base formula?

terenzreignz (terenzreignz):

Also, are you sure it's not like THIS: \[\Large e^{2x}= e^{x + \frac{\log 5}{\log e}}-6\]

OpenStudy (anonymous):

\[\frac{ 1 }{ \log_{b} a }\] right?

OpenStudy (anonymous):

Maybe.. the TA posted the problems and then sent an email saying there was a typo.

terenzreignz (terenzreignz):

Well, you better confirm, because this can be the difference between a simple or outrageously and impossibly complicated answer.

OpenStudy (anonymous):

he takes forever to respond. I'll copy and paste his correction he sent.

terenzreignz (terenzreignz):

You do that :)

OpenStudy (anonymous):

the exponent should be x + log(5)/log(e), rather than x log(5)/log(e).<< this is what he emailed us.

terenzreignz (terenzreignz):

Haha... my instincts serve me correctly, don't they? :P

OpenStudy (anonymous):

here is the original email

terenzreignz (terenzreignz):

yeah, so it really is \[\huge e^{2x}= e^{x+\frac{\log 5}{\log e}}-6\]

OpenStudy (anonymous):

okay whew... I tried it on mathway the other way and it said no answer.. I was getting aggrivated.

OpenStudy (anonymous):

so how do I even begin this problem.. I've looked through our book and haven't seen anything like it

terenzreignz (terenzreignz):

hang on... lag

terenzreignz (terenzreignz):

First, what is \[\Large \frac{\log 5}{\log e}\]